OpenStudy (anisah.21a):

(1/3)^7 (1/3)^3 (-1/3)^8 simplify this expression

jhonyy9 (jhonyy9):

use 1/x = x^(-1)

OpenStudy (fibonaccichick666):

so if this was \[a^7 *a^3\] what would that simplify to?

OpenStudy (anisah.21a):

a^10

OpenStudy (fibonaccichick666):

alright, good, so for yours to start a=1/3

OpenStudy (fibonaccichick666):

so now for the -1/3

OpenStudy (fibonaccichick666):

let's think for a moment. Exponents are just shorthand for multiplication, right?

OpenStudy (anisah.21a):

right

OpenStudy (fibonaccichick666):

so, what's the difference between 5*5 and -5*-5?

OpenStudy (anisah.21a):

no difference, both equals 25

OpenStudy (fibonaccichick666):

alright, so what about 2*2*2 and -2*-2*-2?

OpenStudy (anisah.21a):

8 and -8

OpenStudy (fibonaccichick666):

alright, so when you have a negative number to an even power, do you think the answer will be positive or negative?

OpenStudy (fibonaccichick666):

how about for an odd power?

OpenStudy (anisah.21a):

negative for even and positive for odd

OpenStudy (fibonaccichick666):

so -5*-5 was negative? that's (-5)^2

OpenStudy (anisah.21a):

no positive...

OpenStudy (anisah.21a):

i dont know where you're going with this lmao

OpenStudy (fibonaccichick666):

well what's gonna be the difference between \[(1/3)^8~~and~~(-1/3)^8\]

OpenStudy (anisah.21a):

same answer

OpenStudy (fibonaccichick666):

right, so you said before if you had a^3 * a^7 you get a^(10) well, so now, you said that there is no difference between the positive 1/3 to the eighth and negative 1/3 to the eighth. that means we essentially have a^3*a^7*a^8

OpenStudy (fibonaccichick666):

which, you can easily simplify :)

OpenStudy (anisah.21a):

1?

OpenStudy (fibonaccichick666):

how did you get 1?

OpenStudy (fibonaccichick666):

what's a^(10)*a^8?

OpenStudy (anisah.21a):

i used 1 for a then plugged it into my calculator

OpenStudy (anisah.21a):

a^18

OpenStudy (fibonaccichick666):

why would you pick an a value? our a is 1/3

OpenStudy (fibonaccichick666):

and a^18 is correct

OpenStudy (anisah.21a):

so 1/3^18

OpenStudy (fibonaccichick666):

have to use exponent rules

OpenStudy (fibonaccichick666):

and yes

OpenStudy (anisah.21a):

thanks for explaining!

OpenStudy (fibonaccichick666):

np, I hope you get the process

OpenStudy (anisah.21a):

i do understand now! don't multiply the fractions together if they are all the same (: thanks.